1. The instructions for an electric lawn mower suggest that an A gauge extension cord ( cross sectional area = 4.2 x 10-7 m2 ) should only be used for distances up to 35 m. The resistivity of copper (used in the extension cord) is 1.72 x 10-8 .m at 20C and the temperature coefficient of resistivity of copper is 0.004041 (C)-1.

What is the resistance of type A extension cord of length 35 m at 40C

2. A charge of -100 μC is placed at one corner of an equilateral triangle with sides 0.1 m. A charge of +200 μC is placed at an adjacent corner. What is the potential at the other corner?

3. The membrane surrounding a living cell consists of an inner wall and an outer wall separated by a small distance. Assume that the membrane acts like a parallel plate capacitor in which the effective charge density on the inner and outer walls has a magnitude of 14.5 x 10-6 C/m2. What is the magnitude of the electric field within the cell membrane?

4. At a distance R from a point charge the magnitude of the electric field created by the charge is 750 N/C; at a closer distance , r, the magnitude of the electric field is 2000 N/C. What is the ratio r/R?

5. In a vacuum two point charges are separated by 0.75 m. Charge 1 has a charge of -5.6 x 10-6 C and experiences an attractive force of 25.2 N. What is the magnitude and sign of the second charge?
Answer #1

1) R = pL/A

R = 1.72*10^-8*35 / 4.2*10^-7

R = 1.43 ohm

R = Ro ( 1+ T)

R = 1.43(1 + 0.004041*(40-20))

= 1.54 ohm

so resistance at 40 o C is 1.54 ohm